HUBBLE SPACE TELESCOPE DAILY REPORT #1146 PERIOD COVERED: 0700 EST Daily Status Report as of 145/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed S/C 4974 (DF-224 FS Installation 9.0 and Higher) The installation of Flight Software 9.3 was executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 5560 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1) The WF/PC-2 was used to make several internal bias, K-Spot, and flat calibration observations to monitor the CCD camera electronics, the optical alignment, and to check for contamination. A total of ten images was taken. There was one case of a bias jump, but no intensity gradients were seen. 1.3 Completed FOS/BL 5418 (The Origin of the Ultraviolet Spectral Components) The Faint Object Spectrograph (Blue side) made observations of STAR-075117+144424 to obtain a rapid mode time resolved spectra. A total of eight observations were taken. Five observations were received and displayed, and they were reported to be okay. Three more observations remain on the Science Tape Recorder, and will be analyzed later. 1.4 Completed WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3: Darks) The WF/PC-2 was used to obtain several 30-minute long, dark exposures to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled. Two observations were received and displayed. They were reported to be nominal. More data remains on Science Tape Recorder #1, and will be analyzed later. 1.5 Completed WF/PC-2 5369 (HST Medium-Deep Survey: Cycle 4 Medium, High Galactic Latitude) The WF/PC-2 made a series of observations of high latitude objects as part of a medium deep sky survey. The observations were taken in parallel with other prime science observations. The observations were executed as scheduled. One observations was received and displayed. It was reported to look fine. More data still remains on Science Tape Recorder #1, and will be analyzed later. No anomalies were noted. 1.6 Completed WF/PC-2 5561 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 2: Flats) The WF/PC-2 was used to make a visual flat calibration image to monitor flat field stability and the stability of the analog-to-digital converters. The observations were executed as scheduled. The data still remains on Science tape Recorder #1, and will be analyzed later. 1.7 Completed FOS/BL 5539 (Cycle 4 FOS Photometric Monitor) The Faint Object Spectrograph (Blue Side) used standard star BD+28D4211 to calibrate the absolute sensitivity of its detectors. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 2 Scheduled Reacquisitions : 12 Successful : 12 2.2 FHST Updates: Scheduled :3 Successful :3 2.3 Significant HSTARs: None 2.4 Operations Notes: An anomaly-free operational day was experienced by all HST subsystems. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Three Sets of WF/PC-2 5561 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 2: Flats) 3.2 Complete Two Sets of WF/PC-2 5092 (GTO Low Latitude Parallel Program Cycle 4) 3.3 Complete GHRS/2 5625 (GHRS Side 2 Carrousel Functional Test: Cycle 4) 3.4 Complete WF/PC-2 5450 (Host Galaxies and Cluster Environments of Quasars) 3.5 Complete GHRS/2 5549 (GHRS Side 2 Dark Noise Monitor: Cycle 4) 3.6 Complete WF/PC-2 5397 (Determination of the Extragalactic Distance Scale: Cycle 4 ) 3.7 Complete WF/PC-2 5308 (High Resolution PC Imaging of a Z=2.4 Galaxy: Cycle 4 Medium) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-